I have a model from Gretl which I'm trying to copy the latex code for so l can paste in a web browser. However, when l submit the code I'm getting weird results and the table is not coming out how l would like. How do l get the latex code from the Gretl model to format exactly the same on this web browser.一 回 X gretl: model 1 启 File Edit Iests ave Graphs Analysis LaTex Model 1: OLS, using observations 1-60 Dependent variable: l

Latex Code from copying the output:

\begin{center}

Model 1: OLS, using observations 1--60\\
Dependent variable: l\_Y\\

\vspace{1em}

\begin{tabular}{lr@{.}lr@{.}lr@{.}lr@{.}l}
&
\multicolumn{2}{c}{Coefficient} &
\multicolumn{2}{c}{Std.\ Error} &
\multicolumn{2}{c}{$t$-ratio} &
\multicolumn{2}{c}{p-value} \\[1ex]
const &
17&9965 &
1&20719 &
14&91 &
0&0000 \\
X1 &
$-$0&136205 &
0&0240726 &
$-$5&658 &
0&0000 \\
l\_X2 &
$-$5&19821 &
0&419033 &
$-$12&41 &
0&0000 \\
\end{tabular}

\vspace{1ex}
\begin{tabular}{lrlr}
Mean dependent var & 1.828866 & S.D. dependent var & 0.480144 \\
Sum squared resid & 3.265365 & S.E. of regression & 0.239347 \\
$R^2$ & 0.759930 & Adjusted $R^2$ & 0.751507 \\
$F(2, 57)$ & 90.21549 & P-value($F$) & 2.19\textrm{e--18} \\
Log-likelihood & 2.192880 & Akaike criterion & 1.614240 \\
Schwarz criterion & 7.897273 & Hannan--Quinn & 4.071879 \\
\end{tabular}


\end{center}
물| Paragraph ▼ B 1 := E styles Path: p span

一 回 X gretl: model 1 启 File Edit Iests ave Graphs Analysis LaTex Model 1: OLS, using observations 1-60 Dependent variable: l i coefficient std. error t-ratio p-value 17.9965 1.20719 14.91 2.52e-021 const -0.136205 0.0240726 -5.658 5.19e-07* -5.19821 -12.41 1 X2 0.419033 7.85e-018 ** Mean dependent var 1.828866 S.D. dependent var 0.480144 Sum squared resid 3.265365 .E. of regression 0.239347 0.759930 Adjusted R-squared 0.751507 90.21549 P-value (F) R-squared F (2, 57) 2.19e-18 Log-likelihood 2.192880 Akaike criterion 1.614240 4.071879 Schwarz criterion 7.897273 Hannan-Quinn Log-likelihood for Y =-107.539
물| Paragraph ▼ B 1 := E styles Path: p span
Answer #1

\documentclass[11pt, twoside]{article}
\usepackage{setspace}
\doublespacing
\begin{document}
\noindent Model 1: OLS, using observations 1--60\\
\noindent Dependent variable: ${\rm l}\_{\rm Y}$\\
\vspace{1em}
\begin{tabular}{lr@{.}lr@{.}lr@{.}lr@{.}l}
&
\multicolumn{2}{c}{Coefficient} &
\multicolumn{2}{c}{Std.\ Error} &
\multicolumn{2}{c}{$t$-ratio} &
\multicolumn{2}{c}{p-value} \\
\cline{1-9}
const &
17&9965 &
1&20719 &
14&91 &
0&0000$^{***}$ \\
X1 &
$-$0&136205 &
0&0240726 &
$-$5&658 &
0&0000$^{***}$ \\
l$\_$X2 &
$-$5&19821 &
0&419033 &
$-$12&41 &
0&0000$^{***}$ \\
\end{tabular}\\
\begin{tabular}{lrlr}
Mean dependent var & 1.828866 & S.D. dependent var & 0.480144 \\
Sum squared resid & 3.265365 & S.E. of regression & 0.239347 \\
$R^2$ & 0.759930 & Adjusted $R^2$ & 0.751507 \\
$F(2, 57)$ & 90.21549 & P-value($F$) & 2.19\textrm{e--18} \\
Log-likelihood & 2.192880 & Akaike criterion & 1.614240 \\
Schwarz criterion & 7.897273 & Hannan--Quinn & 4.071879 \\
\end{tabular}\\\vskip.1cm
\noindent Log-likelihood for Y=-107.539
\end{document}